### v2

# LogPrism: Unifying Structure and Variable Encoding for Effective Log Compression

Abstract
[…]
I Introduction
Software systems generate logs to record runtime events, errors, and operational states, which are indispensable for system maintenance [16, 4, 10, 34, 39, 40, 28, 43] and performance optimization [15, 3, 42]. However, the sheer volume of log data poses significant storage challenges [9, 5]. Modern large-scale systems can produce terabytes or even petabytes of logs daily [29, 6, 35]. Thus, efficient log compression has become a critical concern for managing storage costs while preserving the analytical value of historical log data [29, 31].
Logs typically follow a “template+variables” pattern, where static strings are interleaved with dynamic runtime parameters. General-purpose compression algorithms like gzip [8] and LZMA [45] fail to exploit the specific characteristics of log data, often resulting in suboptimal compression ratios [36]. To address this, recent research has focused on log-specific compression methods that leverage the inherent semi-structured nature of logs. By separating the template and variable components, parser-based compressors can replace repetitive templates with compact identifiers and compress variable streams with specialized encoding techniques, achieving significantly higher efficiency.
Despite these advancements, existing approaches face a fundamental limitation, i.e., the decoupling of log parsing [44, 18] and compression. Current workflows treat parsers as a black-box pre-processor, which aim to maximize semantic accuracy (i.e., identifying the correct templates) without regard for downstream compression effectiveness. To quantify the impact of this misalignment, we conduct a comprehensive empirical study evaluating four state-of-the-art compressors and nine parsers. Our findings reveal that semantically accurate parsers may produce templates that undermine compression performance, such as over-generalized templates that offload excessive entropy to the variable stream, or over-fitted templates that incur substantial dictionary overhead. Furthermore, the conventional “parse-then-compress” pipeline creates a boundary between the log structure and parameters, preventing the exploitation of deeper redundancies. Specifically, it ignores template-variable correlations, where specific variable values are strongly tied to a template and could be encoded as part of the structure, and inter-variable correlations, where variables within a single log entry co-occur in predictable patterns. By treating these components in isolation, existing methods fail to mine and encode these high-value aggregate patterns.
To overcome these limitations, we propose LogPrism, a log compression framework that unifies structural extraction and variable encoding. Instead of relying on a pre-defined parsing stage, LogPrism constructs a Unified Redundancy Tree (URT) that dynamically models both log structure and variable correlations in an integrated representation. Our approach employs a hierarchical redundancy mining strategy that progressively distills log data through three stages. First, we extract stable log tokens to construct a structural tree, establishing a compact skeleton for subsequent analysis. Second, we extend the skeleton by building variable subtrees to mine frequent “structure+variable” co-occurrence patterns, effectively bridging the gap between templates and parameters. Finally, we execute residual data processing to efficiently handle the remaining high-entropy “long-tail” variables using a specialized sorting and stream normalization pipeline.
[…]
In summary, this paper makes the following contributions:
  • We conduct the first comprehensive empirical study to quantify the impact of log parsers on compression, revealing the critical misalignment between parsing accuracy and compression efficiency.
  • We propose the concept of unified redundancy encoding, a paradigm shift that co-designs structural extraction and variable encoding to exploit deep “structure+variable” correlations. Based on this concept, we design and implement LogPrism, a high-performance log compression framework featuring a unified redundancy tree and a parallel-aware architecture.
  • We perform an extensive evaluation demonstrating that LogPrism significantly outperforms existing state-of-the-art methods in both compression ratio and speed.
II Background and Motivation
II-A Parser-based Log Compression
In the realm of log compression, parser-based methods represent the predominant approach. They leverage the inherent semi-structured nature of logs, typically composed of static format strings and dynamic parameters. The process begins with a critical prerequisite step, i.e., log parsing. As illustrated in Fig. 1, parsers like Drain [14] are employed to decompose raw log messages, separating structured headers (e.g., timestamps, log levels) from the free-form body, and then segment the body into an invariant event template and its corresponding variables. This structured representation enables the core compression mechanism, where repetitive template text is replaced with compact identifiers (stored only once in a dictionary) and the extracted parameters are aggregated for subsequent encoding using specialized compression techniques. The resulting output, which is a highly regular stream of template IDs and parameter arrays with significantly reduced entropy, is then processed by general-purpose algorithms like LZMA or gzip to eliminate any remaining statistical redundancy. Several representative log compressors [23, 33, 21, 38] exemplify different design philosophies within this framework.
Figure 1: The General Workflow of Parser-based Log Compression
The performance of these parser-based compressors is fundamentally tied to the quality of the log parsing stage [38, 20]. However, existing research mainly focuses on compression algorithms that operate on the parsed logs. Their designs, therefore, default to an assumption of ideal parsing quality. The parser is often treated as an independent component rather than an integral part of the compression pipeline whose performance demands critical evaluation. This has led to a common practice where an off-the-shelf parser is selected based on secondary criteria like execution speed or ease of integration. However, there is no one-size-fits-all parser. Log parsers employ a variety of heuristics and algorithms, and their accuracy can vary significantly across different datasets with diverse log formats. Any errors introduced during parsing [19, 18, 22, 13] will inevitably propagate and undermine the effectiveness of the downstream compression algorithms. Despite this critical dependency, the interplay between parsing accuracy and compression efficiency remains largely unexplored. To systematically evaluate the impact of different log parsers on the performance of log compression, and to further reveal the critical role of parsing in the compression pipeline, we conduct a comprehensive empirical study in this paper.
II-B Revisiting Log Parsers in Compression Pipelines
In this section, we perform an empirical study to investigate how the choice of log parsers quantitatively affects the performance of existing log compressors. We collect a set of representative log parsers and parser-based log compressors, and perform controlled experiments across multiple benchmark datasets that are widely used in log analysis domain.
II-B1 Experiment Setup
We introduce the experiment setup of our study as follows:
Log Parser Selection: We select nine representative log parsers, i.e., Drain [14], AEL [17], IPLoM [24], LFA [27], LogSig [30], MoLFI [25], SHISO [26], Spell [11] and the parser implemented in LogReducer [33]. By covering this broad spectrum of algorithmic strategies, our study can generate heterogeneous template sets and provide a robust evaluation of their impact on downstream compression performance. These parsers span seven distinct methodological families, from simple heuristics to complex data-driven models. For instance, AEL represents the heuristic approach that uses lightweight, generic rules to distinguish static text from dynamic parameters, requiring minimal domain knowledge. In contrast, Drain and LogReducer employ fixed-depth parsing trees, where root-to-leaf paths define templates, enabling efficient online processing. Other parsers frame the task as a data mining problem. LFA applies frequent pattern mining and IPLoM iteratively partitions log messages based on token position and cardinality. Clustering-based parsers like LogSig and SHISO identify templates by computing pairwise message similarity and treat each resulting cluster’s centroid as a template. Our selection also includes other novel strategies. Spell identifies templates based on the longest common subsequence algorithm, making it particularly effective for streaming data. MoLFI leverages evolutionary algorithms to iteratively evolve optimal template sets via genetic operations.
Log Compressor Selection: To select representative log compressors, we review existing solutions [23, 5, 33, 37, 21, 38, 1, 29, 32] in the literature and apply two filtering criteria. First, we exclude online compressors [5, 37]. Our study focuses on offline lossless compression for long-term storage, which prioritizes maximizing storage efficiency across the entire dataset. In contrast, online methods optimize for reducing real-time transmission overhead, often operating under tight resource constraints and with the possibility of lossy compression [41, 37, 2]. Second, we prioritize selecting the state-of-the-art approaches that represent the most effective methodologies in the field. This process yields four representative compressors: Logzip [23], LogReducer [33], LogShrink [21], and Denum [38]. While Denum is primarily a number-centric approach designed to bypass traditional template extraction, we include it because it relies on parser-based strategies like LogShrink to compress its non-numeric log content. Consequently, how the input text is parsed or tokenized remains a critical factor in Denum’s overall performance.
These approaches feature a different design in log structure extraction and encoding strategies, providing a diverse testbed for our study. Logzip, as a pioneer in parser-based compression, employs iterative clustering to uncover latent structures, transforming logs into a hierarchical representation with template identification and parameter mapping. LogReducer advances this idea by focusing on inter-parameter correlations, utilizing an elastic numeric encoding scheme to dynamically select compact bit representations, complemented by delta encoding for sequential data. Similarly, LogShrink optimizes parameter storage by reorganizing data into columnar formats, grouping homogeneous types to create low-entropy streams ideal for dictionary encoding [46]. Finally, Denum introduces a number-centric approach that targets numerical data, using regular expressions to isolate numeric strings for categorization and differential encoding to minimize redundancy.
TABLE I: Detailed Statistics of Benchmark Log Datasets
System Type Dataset File Size # Lines
Distributed Systems HDFS 1.47 GB 11,175,629
Hadoop 48.61 MB 394,308
Spark 2.75 GB 33,236,604
Zookeeper 9.95 MB 74,380
OpenStack 58.61 MB 207,820
Supercomputers BGL 708.76 MB 4,747,963
HPC 32.00 MB 433,489
Thunderbird 29.60 GB 211,212,192
Operating Systems Windows 26.09 GB 114,608,388
Linux 2.25 MB 25,567
Mac 16.09 MB 117,283
Mobile Systems Android 183.37 MB 1,555,005
HealthApp 22.44 MB 253,395
Server Applications Apache 4.90 MB 56,481
OpenSSH 70.02 MB 655,146
Standalone Software Proxifier 2.42 MB 21,329
Dataset Selection: To ensure generalizability and robustness, our evaluation utilizes the widely recognized LogHub benchmark [44]. As detailed in Table I, this collection comprises 16 datasets spanning a broad spectrum of system types, including large-scale distributed systems (e.g., HDFS, Spark), supercomputers (BGL, Thunderbird), server applications (Apache, OpenSSH), operating systems (e.g., Linux, Windows), mobile platforms (Android, HealthApp), and standalone software (Proxifier). In total, the full benchmark contains over 77 GB of raw logs and approximately 378 million log entries. However, our preliminary experiments revealed that the Android and Windows datasets could not be processed by some parsers within a reasonable time budget due to their inherent complexity. Consequently, we exclude these two datasets from our empirical study. Our final evaluation is conducted on the remaining 14 datasets, which collectively account for over 50 GB of data and 262 million log entries, spanning diverse log formats and structural complexities.
Evaluation Metrics: To quantify how different log parsers impact the effectiveness of downstream compression, we employ the Compression Ratio (CR), a standard metric in data compression. It is defined as the ratio of the original log file size to the compressed file size. A higher value indicates better compression performance.
$Compression~Ratio~(CR)=\frac{Original~Log~Size}{Compressed~File~Size}$
II-B2 Experiment Workflow
We design a controlled workflow for the study, which is structured into three phases, i.e., log parsing, intermediate normalization, and log compression.
In the initial phase, we apply the nine selected log parsers to generate structured template collections for each dataset. During this process, we observe significant inconsistencies in parser outputs regarding wildcard notation, delimiter usage, and file formats. For example, some parsers use <*> to denote variables, while others use markers like spec or simply omit the variable tags. Such heterogeneity creates format incompatibilities that can prevent downstream compressors from correctly recognizing templates. To eliminate these discrepancies, we perform intermediate normalization in the second phase. This standardizes all parser outputs into a single canonical format, while preserving semantic content. Specifically, we unify all wildcard symbols, align metadata structures to a uniform schema, and convert outputs to a consistent file format. This ensures that compression performance differences reflect parsing quality rather than formatting artifacts. In the final phase, we evaluate the four selected compressors using these normalized templates. This presents a technical challenge, as most off-the-shelf log compressors are monolithic systems with tightly coupled parsing and compression modules. To address this, we systematically refactor their source code, decoupling data ingestion from the core compression logic. We develop dedicated input interfaces that allow us to inject our normalized templates, creating a controlled environment to test each compressor against templates from all nine parsers.
II-B3 Experimental Results
Fig. 2 presents the compression ratio achieved by four log compressors when supplied with templates generated by nine different parsers. Our analysis reveals three key observations that fundamentally challenge the current parser-based paradigm for log compression.
Figure 2: Impact of Log Parser Selection on Downstream Compression Ratio across Different Datasets
Observation 1: Dramatic Performance Variance Induced by Parser Selection. For any given compressor, the choice of parser induces dramatic variance in the final compression ratio, often surpassing the inherent performance differences between the compressors themselves. Three compressors (LogZip, LogReducer, and LogShrink) exhibit particularly strong sensitivity. For example, on Zookeeper, LogShrink achieves a CR of 121.06 when using templates from LogReducer’s parser, but this plummets to just 44.81 with the IPLoM parser’s templates. Denum, which is not fully parser-based, also demonstrate notable performance fluctuations, confirming that the handling of non-numeric text remains critical. Such results provide clear evidence that parser selection plays a critical, yet previously overlooked, role in compression efficiency.
Observation 2: Default parsers are not always optimal. Counter-intuitively, our experiments reveal that compressors can achieve higher compression ratios using external parsers rather than their built-in ones. On BGL, for instance, LogReducer achieves a CR of 38.04 when using its own default parser, but this increases to 43.08 when paired with the SHISO parser. A similar trend is observed on the HealthApp dataset, where SHISO again outperforms LogReducer’s native parsing logic (29.98 vs. 27.71). These findings suggest that the tight coupling of specific parsers with compressors in current designs can be suboptimal, preventing the compression algorithms from realizing their full potential on diverse datasets.
Observation 3: Parsing Accuracy Does Not Guarantee Compression Efficiency. Our results demonstrate that there is no universally optimal parser capable of consistently maximizing compression efficiency across all datasets and compressors. While Drain generally performs well, it is significantly outperformed by SHISO and LFA on Thunderbird. Conversely, SHISO excels on BGL but yields suboptimal results on HDFS and OpenStack. Even the ground-truth templates (labeled “Correct”) are frequently outperformed by heuristic parsers. This variance indicates that a parser’s effectiveness for compression is not intrinsic but depends on the complex interplay between data characteristics and the compressor’s encoding strategy. The root cause of this volatility lies in a fundamental misalignment of objectives: parsers prioritize classification accuracy to maximize event clustering correctness, whereas compressors prioritize storage efficiency. Parsers emphasize semantic accuracy without considering the storage cost of the resulting templates and parameters. This leads to two typical inefficiencies that degrade compression performance despite high parsing accuracy:
  • Over-generalization (Coarse-grained Templates): Parsers prioritizing high matching rates with fewer templates often introduce excessive wildcards. On HDFS, for instance, LFA generates only 42 templates but with a total of 372 wildcards (8.8 per template on average). While this may yield high parsing accuracy, it offloads complexity to the parameter stream. The compressor is forced to process large, noisy parameter sets that should have been static template text, severely degrading the compression ratio.
  • Over-fitting (Fine-grained Templates): Conversely, overly sensitive clustering thresholds can partition semantically similar logs into numerous distinct templates. For example, some parsers produce over 28,000 templates for the HealthApp dataset, which contains only around 150 actual event types. This “template explosion” exponentially increases the dictionary size, creating a direct storage overhead that undermines compression efficiency despite high formal parsing precision.
In summary, our empirical study confirms that log parsing is not a mere preprocessing step but a dominant factor in the ultimate efficiency of log compression. The fundamental principle of exploiting log structural redundancy is sound, which enables parser-based approaches consistently and significantly outperforming general-purpose algorithms like gzip or LZMA. However, our findings reveal a critical flaw in its common decoupled implementation. We argue that unlocking optimal log storage performance requires a paradigm shift where structural extraction is not a prerequisite for compression but an integral, co-designed part of it.
Figure 3: A 12-line Log Snippet as a Running Example
II-C A Motivating Example for Unified Redundancy Encoding
In this section, we illustrate the core concept of our approach based on the logs in Fig. 3. The idea is to move beyond the rigid dichotomy of “static strings vs. dynamic variables” by modeling log tokens (whether traditionally considered a static string or a variable) uniformly and encode them in aggregate based on their co-occurrence and dependencies. To this end, we propose a paradigm shift from the traditional parse-then-compress workflow and introduce a new methodology called unified redundancy encoding.
Consider the sshd log entries in Fig. 3 (L5-L7 and L9-L12). Since variable user, rhost, and uid take different values, a conventional parser may generate a log template like "...authentication failure; user= <*> rhost=<*> uid=<*> euid=0". When compressing these logs, e.g., L5, existing compressors would first store an identifier for this template and then independently encode the variable values (test1, pokemon1.cs.edu, and 509). In this process, two types of correlations are ignored. The first is template-variable correlation. For pure log compression purposes, the template ID could include certain variables if they are strongly tied to the template body, eliminating separate parameter processing. Although parsers may occasionally inline frequent variables (e.g., euid=0 in Fig. 3) into the template, this behavior is inconsistent and not measurable. The second is inter-variable correlation within each log entry, which allows for the collective processing of variables. For instance, user test1 deterministically maps to uid 509, and user root maps to uid 0. This differs from prior correlation mining methods [21, 38] that mainly model relationships across log entries (e.g., incremental user IDs).
Our method is more holistic and context-aware. It recognizes that the value set $\{\texttt{test1},\texttt{pokemon1.cs.edu},\texttt{509}\}$ (L5 and L7) is a frequent pattern occurring within the context of the aforementioned template. Therefore, our method aggregates both parts as "...authentication fail- ure; user=test1 rhost=pokemon1.cs.edu uid= 509 euid=0" and encodes it with a single ID. Similar frequent patterns include $\{\texttt{root},\texttt{pc180.edu.tw},\texttt{0}\}$ (L9 and L11) and $\{\texttt{root},\texttt{julia.arkos.de},\texttt{0}\}$ (L10 and L12), while only srv2.alfahost.nl (L6) needs to be handled independently. The advantage of this unified redundancy encoding is significant: whereas existing approaches require storing one template ID plus three separate variable IDs, we need only a single ID to represent the entire, highly correlated token sequence. This principle of co-designing structural extraction and pattern encoding enables us to discover and exploit deep contextual redundancies in log data.
A straightforward way to implement this idea is to build a prefix tree (i.e., a trie) over the sequences of log tokens and assign IDs to frequent paths, thereby identifying recurring log (sub)sequences. However, this naive construction presents two fundamental limitations. First, high-cardinality fields induce many branches, causing the number of nodes to grow with the product of distinct values per position. This results in prohibitive storage consumption and increased lookup latency due to structural inefficiency. Second, a trie only encodes prefixes. If an infrequent token appears early, the match stops and the rest of the log line is no longer compressible, even when the suffix is highly regular (e.g., uid=509 euid=0 after a rare rhost in L6). These constraints motivate our design of LogPrism, an effective log compressor which moves beyond pure prefix matching and supports compact, gap-tolerant aggregation of frequent token sequences.
III Methodology
Figure 4: An Overview of the LogPrism Compression Framework
In this section, we introduce LogPrism, our log compression framework that constructs a Unified Redundancy Tree (URT) to jointly model log structure and variable correlations. Since a naive prefix tree over raw logs would suffer from combinatorial branch explosion, LogPrism employs a hierarchical redundancy mining strategy to build a structurally efficient URT. We progressively integrate log tokens into the URT based on their frequency and stability through three stages, i.e., Structural Tree Construction, Variable Subtree Encoding, and Residual Data Processing, as shown in Fig. 4. The first stage builds the structural skeleton of the URT based on regular, fixed tokens. Each path in this tree represents a group of log entries that are structurally similar, regardless of the log events they represent. The second stage integrates frequent variable tokens by expanding the skeleton’s terminal nodes into subtrees. Within each subtree, we mine deep correlations between the structural skeleton and specific variable values, allowing a single ID to represent complex, recurring “structure-variable” patterns. The third stage handles the remaining “long-tail” tokens (i.e., outliers that are too rare or random) by isolating and encoding them via specialized schemes suited for high-entropy data. The entire compression pipeline concludes by feeding all compressed data into a general-purpose compressor (e.g., LZMA) to exploit remaining byte-level redundancy.
It is essential to note that LogPrism’s hierarchical compression pipeline is fundamentally different from the decoupled “parse-then-compress” workflow. LogPrism’s ultimate goal is to unify the encoding of both structural and variable tokens into an end-to-end redundancy pattern, rather than treating them as separate entities for isolated compression.
III-A Structural Tree Construction
The primary objective of this stage is to construct the foundational skeleton of the URT by extracting stable patterns from the raw logs. To ensure high throughput and a low memory footprint, we implement this process within a parallel streaming architecture. The log dataset is partitioned into multiple chunks, each processed concurrently by a dedicated worker thread to build local structures. As depicted in Fig. 4, this construction phase proceeds through two key steps: 1) Parallel Pre-processing and Local Tree Generation, where logs are tokenized, filtered, and organized into local prefix trees; and 2) Global Tree Aggregation and Isomorphic Subtree Merging, which integrates these local trees into a unified global structure while dynamically refining the topology.
III-A1 Parallel Pre-processing and Local Tree Generation
The pipeline within each worker thread involves two operations, i.e., the pre-processing of the assigned log chunk to filter volatile content, and the construction of a local structural tree.
To prevent the “branch explosion” associated with high-cardinality data, we identify and separate two categories of rapidly-changing tokens.
  • Globally Patterned Metadata: This category includes common header fields like dates, timestamps, and process IDs (PIDs) [44]. These tokens, while highly variable, exhibit predictable syntactical patterns. Storing them in a string-based tree is highly inefficient. Thus, we identify them using predefined regular expressions and extract their values into separate, highly compressible columnar streams, preserving the global order of the logs. As illustrated in Fig. 5(a), these tokens are replaced with specific placeholders (e.g., $\langle X\rangle$ for month, $\langle dt\rangle$ for timestamp, $\langle P\rangle$ for PID) in raw logs.
  • Unstructured Numeric Tokens: This category targets tokens within free-text log contents that are likely to be variables. Based on a simple and effective heuristic from prior research [38], we treat tokens containing numeric characters as potential variables. We apply this rule to tokens not captured by the metadata regex and uniformly replace the matched ones by the wildcard placeholder $\langle*\rangle$, as shown in Fig. 5. The original values of these tokens are collected sequentially into a varList associated with the specific log entry. In this process, two types of misclassifications can happen. First, static tokens containing numbers (e.g., node1) are generalized as variables. Since LogPrism treats static and dynamic tokens as a holistic entity, these tokens can be re-integrated into the structure during correlation mining stage (Sec. III-B). Second, string-only variables (e.g., user=root) are treated as static, causing potential tree fragmentation. This is explicitly addressed via isomorphic subtree merging in the subsequent step.
After pre-processing, the resulting log messages are used to construct a local prefix tree. For example, dividing the logs from Fig. 3 into two chunks (L1-L6 and L7-L12) yields the independent trees shown in Fig. 6(a) and Fig. 6(b), respectively. In this structure, every terminal node marks the end of a log message. Crucially, a terminal node is not necessarily a leaf node, since one log entry can be a complete prefix of another. Each path from the root to a terminal node uniquely identifies a group of log entries sharing the same structural pattern. To index these groups, a Record Object is maintained at every such terminal node, aggregating the Line IDs (Lid) and extracted varList(s) for all corresponding logs.
Figure 5: The Pre-processing Pipeline for Log Entry L5
Figure 6: The Parallel Pre-processing and Streaming Merge Process of the Structural Tree
III-A2 Global Tree Aggregation and Isomorphic Subtree Merging
Once the worker threads generate their local structures, the main thread integrates them into a single global tree and refines the topology. We employ an on-arrival merge strategy to minimize synchronization overhead. As soon as a worker thread completes its chunk, its local prefix tree is merged into the global one. As depicted in Fig. 6(c), this merge process aggregates information along identical paths and adds new branches where structures differ (e.g., the user=root branch from the second chunk). Concurrently, the columnar data streams from each thread are appended to global files, preserving the original order of the log dataset. For numerical streams like $\langle dt\rangle$, we apply a specialized compression pipeline consisting of Delta Encoding (to store value differences), ZigZag Encoding (to efficiently represent negative numbers), and Varint Encoding (for variable-length integer representation) to generate a final compact binary format.
Although the number-based heuristic for identifying volatile fields in the first step is effective, it can potentially misclassify purely string-based variables. For instance, user=test1 is correctly generalized to $\langle*\rangle$ (due to "1"), but user=root would be incorrectly identified as a stable token. This inconsistency creates unnecessary branches and fragments the tree. To resolve this, we introduce a correction process named isomorphic subtree merging. It operates on the principle that if a “static” branch and a “variable” branch lead to subtrees that are topologically isomorphic, they serve the same structural role and should be treated uniformly as a variable.
The merging process is implemented as a post-order (bottom-up) traversal of the global tree. At each branching node, it computes a unique structural signature for every child subtree. If a node contains a linear path, its signature is simply the concatenation of its tokens. For nodes with multiple branches, we lexicographically sort the signatures of all outgoing branches before concatenating them. Any sibling nodes sharing identical structural signatures will be considered isomorphic, and the algorithm performs a merge operation. Fig. 7 demonstrates this iterative merging. The traversal moves up from the leaves and, as shown in Fig. 7(a), encounters its first branching node user=root. It compares the child nodes $\langle*\rangle$ (which represents numeric rhosts) and rhost=julia.arkos.de by computing the structural signature for the path under each. Since both yield the same sequence of $\langle*\rangle$ nodes (as highlighted by the green dashed box), they are topologically identical. This isomorphism indicates that rhost=julia.arkos.de should also be a variable token (i.e., $\langle*\rangle$). The system then merges these two paths by (i) inserting the value "rhost=julia.arkos.de" into the corresponding position of the varList (the red arrow line) for all logs traversing that path, and (ii) merging the Record Object (i.e., Lids and updated varLists) from the string branch into the sibling wildcard branch. The result is the more generalized structure in Fig. 7(b). As the traversal continues up to the next branching node, failure;, the system performs another similar merge as shown in Fig. 7(c).
The choice of a bottom-up traversal is critical for accurate path mergings. By processing the tree from the leaves upward, the algorithm ensures that sibling nodes sharing the longest common prefixes are evaluated for merging first. Such paths are more likely to be truly isomorphic [14]. It also ensures that any structural inconsistencies at deeper levels are resolved first, recursively generalizing the tree into its most compact form. In case of incorrect mergings, LogPrism can separate the outlier variables during the correlation mining stage (Sec. III-B).
Figure 7: The Iterative, Bottom-up process of Isomorphic Subtree Merging
Finally, we traverse the global tree to index the structural contexts. Every terminal node is assigned a globally unique pathID. As shown in Fig. 8, this ID acts as a structural context identifier, representing a group of logs that share the exact same skeletal structure. For example, all sshd logs (L5-L7 and L9-L12) in the example are mapped to pathID=3. The core role of the pathID is to provide clear log groupings for the next stage of variable correlation analysis, where each log is represented by its original line number (Lid), its structural context identifier (pathID), and its varList.
Figure 8: The Final Global Structural Tree with pathID Assignments
III-B Variable Subtree Encoding
The construction of the structural context tree leverages only stable tokens, resulting in a highly compact skeleton where all variable tokens are aggregated at the terminal nodes. To capture the deep correlations between the log structure and its parameters, we can conceptually extend the URT by attaching a prefix subtree to each terminal node, using the varLists of the associated log entries as input paths. By doing so, we unify the structural skeleton and variables into a single, continuous representation. However, this presents two critical challenges. First, a naive sequential insertion of variables inevitably triggers a secondary branch explosion. As illustrated in Fig. 9(a), if variables are processed in their original order (e.g., $v_{0}\to v_{1}\to v_{2}\to v_{3}$), a high-cardinality variable may appear early in the sequence (e.g., a rhost at $v_{1}$), forcing the subtree to branch at the top layers. Consequently, even if later variables form a highly frequent pattern (e.g., uid=509 $\to$ euid=0), it is fragmented into multiple separate branches, preventing efficient compression. Second, we require a robust mechanism to distinguish meaningful patterns from random noise. Since variables are high-entropy, assigning a unique ID to every combination would result in a prohibitively large dictionary and compromise the compression benefits. Therefore, we must selectively encode only frequent, co-occurring variable combinations (the “signal”) into the URT, while actively filtering out rare or random values (the “noise”) to be handled as residuals (Sec. III-C).
III-B1 Variable Reordering for Efficient Subtree Construction
To mitigate the secondary branch explosion, LogPrism restructures the variable subtree topology via a two-step optimization process. This involves a stability-based sort that prioritizes low-entropy variables for insertion. By forcing frequent patterns to share a common prefix, this strategy effectively pushes high-cardinality branching to the deeper levels of the tree. Consequently, correlated variable sequences are materialized as deep, continuous paths (Fig. 8) rather than fragmenting prematurely.
The optimization begins with high-frequency filtering, which prunes infrequent variables, as shown in Fig. 10(a). For each variable position within a log group, we compute the frequency of every unique value and discard any that fall below a pre-defined threshold $\tau$, e.g., the rare rhost=srv2.alfahost.nl. If all values at a given position are filtered out, the entire position is excluded from subsequent analysis. For instance, in the httpd log group (pathID=1) shown in Fig. 8, the sole variable $v_{0}$ (transaction_id) is removed entirely because both of its values (5001 and 5002) are too infrequent. In this work, we use $\tau$ as our universal threshold for different filtering operations to screen out outliers, which avoids the overhead of managing multiple distinct thresholds and reduces overall system complexity.
The second step is variable reordering, which establishes the optimal insertion order for the remaining variables. The reordering is determined by a heuristic that prioritizes stability. We calculate two key metrics for each variable position, i.e., total frequency (the sum of occurrences of all its values) and discriminative power (the number of its unique values). Variables are then sorted in descending order of total frequency. Any ties are resolved by sorting in ascending order of discriminative power, which favors variables with fewer unique values. In the sshd example, after filtering, $v_{1}$ has the lowest total frequency, i.e., six, while the other positions share a total frequency of seven. Among those, $v_{3}$ has the lowest discriminative power of one. This heuristic thus yields the optimal construction order of $v_{3}\to v_{0}\to v_{2}\to v_{1}$, which is the crucial input for the final subtree construction phase.
Figure 9: The Motivation and Effect of Variable Reordering
Figure 10: The Variable Subtree Encoding Pipeline for the sshd Log Group (pathID=3)
III-B2 Subtree Construction and Unified Redundancy Encoding
With the established variable order, this phase constructs the variable subtrees to identify and encode co-occurrence patterns. While the reordering was guided by aggregate statistics, it does not guarantee that individually frequent variables appear together in the same log entry. For instance, a specific user and rhost might both be globally frequent, but if they never co-occur (e.g., that user never logs in from that host), they do not form a valid pattern. Therefore, this stage validates these patterns by iterating through each log’s varList, ensuring that compression identifiers are assigned only to variable combinations that genuinely exist in the data.
The construction process is illustrated in Fig. 10(b). As each log’s varList is traversed according to the optimal order (i.e., $v_{3}\to v_{0}\to v_{2}\to v_{1}$), a path is extended in the subtree. Every node in this tree maintains a cnt attribute, a key metric that records the number of logs whose reordered variables share that specific prefix. For logs like L5 and L7, which are composed entirely of frequent variables, a complete path is formed, and the cnt value of every node along that path is incremented. In contrast, the traversal for log L6 terminates prematurely because its variable rhost=srv2.alfahost.nl was filtered out as a low-frequency value. Consequently, L6 only increments the cnt values for its matched prefix (euid=0 $\to$ user=test1 $\to$ uid=509). This distinction is crucial, as the cnt attribute now precisely tracks the frequency of both complete and partial patterns. Next, a pruning operation is performed, which acts as a second layer of filtering, removing paths composed of individually frequent variables but whose combination is rare. Any node whose cnt falls below the pruning threshold $\tau$ is removed. While no nodes are pruned in our simplified example (Fig. 10), this is crucial for eliminating noise in complex datasets.
The most critical step is the assignment of new universal pathIDs. To maximize compression efficiency, identifiers are assigned only to nodes that represent the termination of a high-frequency aggregate pattern. This ensures that every pathID corresponds to a meaningful, recurring token combination, preventing an explosion in dictionary size due to rare partial matches. We term these nodes “stable endpoints” and identify them using the cnt attribute under two conditions. First, any leaf node is inherently a stable endpoint, as it represents the explicit termination of a pattern that has already survived the high-frequency filtering process (i.e., $\tau$). For example, the rhost=pokemon1.cs.edu node in Fig. 10(b) is assigned pathID=5 because it marks the end of the pattern for logs L5 and L7. Second, a non-leaf node is designated a stable endpoint if the difference between its cnt and the sum of its children’s cnt values is not less than the threshold $\tau$. This “residual count” represents the number of log entries whose pattern matches exactly up to this node but does not continue to any of its high-frequency children. By enforcing the threshold, we ensure that we only create a new ID if a substantial number of logs terminate at this specific intermediate point. For instance, the uid=509 node has a cnt of 3, while its only child has a cnt of 2. The residual count of 1 indicates that one log (L6) terminates its match here. If this residual meets the threshold, the node is marked as a stable endpoint and assigned pathID=4, allowing LogPrism to hierarchically encode both complete patterns and frequent sub-patterns.
The final encoded output, shown in Fig. 10(c), is a highly compact representation of the log data. Log messages that fully match a frequent path, such as L7 and L9, are represented by a single pathID (5 and 6, respectively). Logs that only partially match, like L6, are represented by the pathID of their longest matched prefix (pathID=4), while the unmatched token (rhost=srv2.alfahost.nl) is preserved as a residual variable in the log’s varList for processing in the third stage. This process is the core manifestation of our unified redundancy encoding concept. Crucially, the new pathID is a logical extension of the initial structural pathID. It represents the complete “structure + variable” collective pattern. Unlike traditional methods that require one identifier for the template and separate ones for each variable, our approach uses a single pathID to represent the entire high-frequency combination. For fully matched logs, this identifier replaces both the original structure and all of its variables, achieving a significant gain in compression efficiency.
III-C Residual Data Processing
After the first two stages have captured high-frequency collective patterns, the remaining data constitutes the “long-tail” data, i.e., outlier variables characterized by high entropy and weak correlation. This final stage is designed to efficiently compress these residual components by leveraging simpler, linear patterns that may exist. Our approach is guided by the principle of efficient residual handling: after the URT has filtered out complex correlations, this final pipeline can focus on simpler sequential regularities.
III-C1 Global Sorting for Temporal Coherence
The structural grouping in the first stage, while effective for mining template-based redundancy, inevitably fragments the natural time order of logs. This can obscure valuable patterns that span across different structural groups. As shown in Fig. 11(a), logs from different sources (e.g., httpd and database) may share a sequential transaction_id. To uncover these dependencies, the core of this stage is a global sorting pipeline. We aggregate the residual information from all log groups and perform a global re-sort based on the original Line ID (Lid). This operation restores the dataset’s temporal coherence, realigning dispersed entries like L3, L4, and L8 and exposing the incremental numeric sequence (5001, 5001, 5002), which is now highly amenable to delta compression.
III-C2 Just-in-Time Residual Templatization
The residual variables after the second stage are a heterogeneous mix of complex strings, atomic identifiers, and pure numbers. This diversity prevents efficient columnar compression. To resolve this, we employ a just-in-time templatization strategy that iterates through the sorted queue and parses each variable into homogeneous components.
For each residual variable, a regex-based extraction mechanism decomposes it into its invariant static fragments and dynamic numeric parts. We then borrow the core principle from Denum [38] to classify each extracted numeric string based on its intrinsic features (e.g., length, first digit) and generate a corresponding placeholder. This unified process gracefully handles three distinct scenarios:
  • Complex Numeric Variables: For a composite variable like audit(1119799950.864:693295): in log L1, the process generates a generalized template from its static parts and placeholders (e.g., audit($\langle jb\_\rangle$$\langle c\_\rangle$$\langle fg\_\rangle$):) and dispatches each numeric value to its respective columnar stream.
  • Atomic String Variables: For a variable that cannot be further deconstructed, such as rhost=srv2. alfahost.nl from log L6, the entire string is treated as an indivisible atomic template and assigned a unique template ID.
  • Pure Numeric Variables: For simple numbers like 5001, a minimalist placeholder template (e.g., $\langle df\_\rangle$) is generated, and the value is appended to the corresponding numeric stream.
The result of this process is a set of dense, homogeneous columnar streams, as shown in Fig. 11(b). The pathID of each log is written to the main stream, followed by the template IDs for its residual variables in the ResVar streams. Concurrently, all extracted numeric values are written to their specialized Value Streams, which are then compressed using Delta Encoding followed by ZigZag and Varint encoding.
Although this stage borrows the principle of numeric classification from Denum, our hierarchical redundancy mining strategy is fundamentally different. Denum performs an indiscriminate, global numeric/non-numeric split from the outset, which can prematurely destroy valuable contextual correlations. In contrast, LogPrism has already processed the vast majority of high-frequency tokens (including many numerics) via unified redundancy encoding in the first two stages. Consequently, the computationally intensive parsing logic of this stage is applied only to a minimal subset of true residual data. This targeted approach provides a dual advantage: a higher compression ratio, as it preserves the “structure + variable” correlations, and faster compression speed, by confining the expensive variable processing to a much smaller dataset.
The pipeline concludes by aggregating all generated components, i.e., the URT, the residual template dictionary, and the various columnar data streams, into a single compact archive, which is finally processed by a general-purpose compressor (e.g., LZMA) to exploit remaining byte-level redundancy.
Figure 11: The Residual Data Processing Pipeline in Stage 3
III-D Decompressor
The decompression process of LogPrism is the precise inverse of compression, achieving lossless log reconstruction. The process begins by loading the metadata from the compressed archive, including the URT, the residual template dictionary, and all columnar data streams. The core of the reconstruction is a sequential traversal of the main pathID stream, reconstructing logs line by line. Since this stream was written in the order of the original log entries, processing it sequentially naturally restores the original log order.
For each pathID read from the stream, a two-stage reconstruction process is performed. First, the pathID is used to perform a reverse lookup in the URT. By tracing backward from the pathID node to the root, we instantly reconstruct the log’s complete high-frequency “structure + variable” pattern. For a log that was fully matched during compression (e.g., L5), this single lookup is sufficient to recover all its tokens, including both the structural and variable tokens. For a log with residual variables (e.g., L1 or L6), this lookup only recovers the matched portion, and the remaining variables are then reconstructed by reading their template IDs from the ResVar streams and resolving them against the residual template dictionary. At this point, a semi-reconstructed log line is formed, containing all static and high-frequency variable tokens but still holding placeholders for dynamic values (e.g., $\langle dt\rangle$, $\langle P\rangle$, and $\langle jb\_\rangle$). Therefore, the second step performs global placeholder substitution. For each placeholder encountered, the decompressor reads the next available value from the corresponding columnar data stream and substitutes it in place. This multi-layered mechanism guarantees 100% accuracy, as all data streams are consumed in the exact order in which they were generated, ensuring a perfect reconstruction of the original log file.
IV Evaluation
We conduct a comprehensive evaluation to demonstrate the effectiveness and efficiency of LogPrism. Our evaluation is designed to answer four critical research questions:
  • RQ1: How does LogPrism’s compression ratio compare to state-of-the-art log compressors?
  • RQ2: How does LogPrism’s compression speed compare to state-of-the-art log compressors?
  • RQ3: What are the individual performance contributions of LogPrism’s hierarchical stages?
  • RQ4: What is the impact of input data granularity on LogPrism’s performance?
IV-A Experimental Settings
IV-A1 Datasets
Consistent with our previous empirical study (Sec.II-B), we utilize the LogHub benchmark [44] for evaluation. However, while the empirical study involved only 14 datasets, this evaluation incorporates the full suite of 16 datasets. The additional datasets (i.e., Android and Windows) were previously excluded due to the scalability limitations of certain parsers. In this evaluation phase, all selected compressors successfully handle these datasets using their default parsers, allowing for a comprehensive assessment across the entire benchmark.
IV-A2 Evaluation Metrics
We employ two fundamental metrics in the field of log compression: the Compression Ratio (CR) to quantify effectiveness (higher values indicate better compression) and the Compression Speed (CS) to measure efficiency (higher values indicate faster processing). The CR was previously defined in Sec. II. The CS is defined as:
$Compression~Speed~(CS)=\frac{Original~File~Size~(MB)}{Compression~Time~(s)}$
IV-A3 Baselines
We compare LogPrism against comprehensive baselines spanning both state-of-the-art log-specific compressors and well-known general-purpose compressors.
  • Log-Specific Compressors: We evaluate LogPrism against the same four state-of-the-art methods studied in our empirical study (Sec. II-B). These include LogZip [23], the pioneering parser-based method; LogReducer [33] and LogShrink [21], which represent advanced approaches optimized for parameter correlation and variability; and Denum [38], the leading number-centric compressor that operates without traditional parsing. This selection ensures a comprehensive comparison against the best existing techniques across different design paradigms.
  • General-Purpose Compressors: We include four compressors in this category: gzip (a traditional compressor based on the DEFLATE algorithm, known for good speed but moderate compression ratios), bzip2 (which uses the Burrows-Wheeler Transform for better compression at the cost of lower speed), LZMA (a dictionary-based algorithm typically offering the best compression ratios but is relatively slow), and PPMd (which compresses by analyzing character sequences and predicting probabilities, excelling on plain text but degrading significantly on logs with high-entropy numeric variables).
For metrics independent of the experimental environment (CR), we directly cite the results from the original papers of the respective log compressors. For environment-dependent metrics (CS), we re-ran all available open-source tools in our unified environment to ensure fair comparison.
IV-A4 Implementation
All experiments were conducted on a Linux server with an AMD EPYC 9224 CPU (24 cores/48 threads, 3.70 GHz) and 251 GB RAM, running Ubuntu 22.04 LTS (kernel 6.8.0). LogPrism is implemented in C++ and compiled with g++ following the C++20 standard. Regular expression matching relies on the PCRE2 library. Similar to Denum, the PCRE2_CODE_UNIT_WIDTH was set to 8.
To align with the evaluation settings of Denum and related work, all input log data for RQ1-RQ3 experiments were split into 100K-line chunks for parallel compression. Each compressor uses 4 threads for processing each chunk. To ensure fair architectural comparison, our main evaluation of LogPrism constrains the internal operations of each chunk-processing thread to be single-threaded, matching the execution model of the baselines. To demonstrate the full potential of our parallel-aware design, we also evaluate an enhanced configuration, denoted as LogPrism-P, where each of the 4 chunk-processing threads utilizes an internal pool of 4 worker threads. The total time to compress all chunks is recorded. LogPrism employs the ‘tar’ utility to package all generated intermediate files, which are then compressed with ‘lzma’ to form the final archive.
IV-B RQ1: The Compression Ratio of LogPrism
Table II presents the compression ratio comparison across all benchmarks. LogPrism achieves the highest compression ratio on 14 of 16 datasets, establishing a new state-of-the-art in compression effectiveness.
[…]
Comparison with log-specific compressors: LogPrism advances the state-of-the-art among log-specific methods. Compared to LogReducer, CR improvements reach up to 59.26%. Against LogShrink, the maximum gain is 56.63%. Notably, even compared to the current leading method Denum, LogPrism achieves substantial improvements of up to 25.93% on Linux and 18.86% on Thunderbird. The LogZip result for Thunderbird is marked as unavailable (–) because, as reported in the LogShrink paper [21], LogZip failed to complete parsing within one week.
TABLE II: Experimental Results of Compression Ratio
Dataset gzip LZMA bzip2 PPMd LogZip LogReducer LogShrink Denum LogPrism
Android 7.742 18.857 12.787 19.370 25.165 20.776 21.857 32.494 32.822
Apache 21.308 25.186 29.557 31.688 30.375 43.028 55.940 58.517 64.330
BGL 12.927 17.637 15.461 18.927 32.655 38.600 42.385 41.804 47.145
Hadoop 20.485 36.095 32.598 32.110 35.008 52.830 60.091 78.546 79.733
HDFS 10.636 13.559 14.059 19.155 26.666 22.634 27.319 25.670 26.466
HealthApp 10.957 13.431 13.843 15.337 12.632 31.694 39.072 44.472 50.477
HPC 11.263 15.076 12.756 14.822 27.208 32.070 35.878 45.275 45.447
Linux 11.232 16.677 14.695 18.508 23.368 25.213 29.252 30.449 38.343
Mac 11.733 22.159 18.074 28.469 26.306 35.251 39.860 40.789 44.992
OpenSSH 16.828 18.918 22.865 31.977 42.606 86.699 103.175 101.654 113.327
OpenStack 12.158 14.437 15.231 17.429 17.258 16.701 22.157 22.238 23.356
Proxifier 15.716 18.982 23.619 25.489 21.493 25.501 27.029 27.288 29.849
Spark 17.825 19.908 26.497 30.614 20.825 59.470 59.739 59.470 54.813
Thunderbird 16.462 27.309 25.428 33.026 49.185 48.434 63.824 75.862
Windows 17.798 202.568 67.533 61.083 310.596 342.975 456.301 481.350 497.298
Zookeeper 25.979 27.667 36.156 38.931 47.373 94.562 116.981 135.251 143.742
This systematic performance advantage of LogPrism stems from a fundamental architectural shift. Traditional methods (whether the “parse-then-compress” workflows or Denum’s “numeric/non-numeric” split) perform an irreversible token categorization early in the process. This leads to the loss of contextual correlations that spans these category boundaries. In contrast, LogPrism’s unified redundancy encoding paradigm co-designs structural extraction with pattern encoding. This enables representing entire, highly correlated “structure + variable” token sequences with single path identifiers, where other methods require one template ID plus multiple independent variable IDs. Our hierarchical redundancy mining strategy prioritizes encoding high-value aggregate patterns while preserving full context, allowing it to exploit deep redundancies in log data.
On the two datasets where LogPrism does not achieve the highest CR (HDFS and Spark), performance remains highly competitive. On HDFS, the difference from optimal is minimal (3.12%). The larger gap on Spark reflects the characteristics of the dataset, i.e., Spark logs are dominated by simple, repetitive templates with single continuously changing numeric values (e.g., over a million records like "Update row <*>"[38]. In this case, the deep correlation mining of LogPrism’s second stage offers less advantage compared to methods specifically optimized for simple numeric streams. Nevertheless, LogPrism achieves a better performance balance in these cases with significantly higher compression speed, as we demonstrate in RQ2.
Summary for RQ1: LogPrism surpasses existing general-purpose and log-specific compressors in compression ratio on 14 out of 16 datasets, establishing a new state-of-the-art in compression effectiveness.
IV-C RQ2: The Compression Speed of LogPrism
We evaluate LogPrism’s efficiency in two configurations: LogPrism, which restricts internal processing to a single thread per chunk for a fair comparison, and LogPrism-P, which activates our internal fine-grained parallelism. A critical aspect of this benchmark is the timing methodology. We measure the processing time for LogPrism (both configurations) and Denum on a strictly end-to-end basis. In contrast, for parser-based baselines (LogZip, LogReducer, LogShrink), we adhere to the conventions established in their respective papers, which exclude the often time-consuming parsing and template generation phases. Consequently, LogPrism’s performance is achieved under a significantly stricter measurement standard than the parser-based baselines.
[…]
The observed variation in LogPrism’s speed across datasets is primarily driven by dataset size, which dictates the effectiveness of chunk-level parallelization. For example, on the small Linux dataset (25,567 lines), the volume is insufficient to fill even a single 100K-line processing chunk, precluding the standard multi-threaded acceleration used by all compressors. Despite this constraint, the standard LogPrism configuration achieves the best performance (8.55 MB/s) among all baselines. Moreover, LogPrism-P provides substantial additional speedup by leveraging the co-design of hierarchical redundancy mining and internal fine-grained parallelism. This ensures high efficiency even in edge cases. On the Linux dataset, LogPrism-P (10.23 MB/s) is significantly faster than Denum (4.97 MB/s) because its hierarchical strategy minimizes the computational overhead, while the internal worker threads maximize CPU utilization within the single active chunk.
TABLE III: Experimental Results of Compression Speed (MB/s)
Dataset LogZip LogReducer LogShrink Denum LogPrism LogPrism-P
Implementation Python C++ C++ and Python C++ C++ C++
Android 0.068 19.323 4.123 24.380 31.189 36.672
Apache 0.737 2.347 1.537 6.369 12.002 17.551
BGL 0.874 26.738 2.571 23.575 33.234 50.625
Hadoop 0.901 12.882 4.401 24.453 40.539 52.840
HDFS 0.701 23.598 3.466 25.193 29.281 35.288
HealthApp 0.736 7.937 2.754 17.866 22.239 29.103
HPC 0.644 9.391 3.599 25.216 32.257 37.868
Linux 0.687 1.249 0.941 4.969 8.550 10.228
Mac 0.009 5.450 2.141 6.710 11.715 16.097
OpenSSH 0.715 14.773 3.335 25.572 51.369 70.016
OpenStack 0.537 13.039 4.018 12.352 20.960 29.396
Proxifier 0.716 1.328 0.742 6.000 9.657 10.585
Spark 0.550 21.233 3.185 26.034 42.052 59.005
Thunderbird 18.656 4.036 19.830 38.539 69.213
Windows 1.357 18.483 6.330 28.783 79.220 116.152
Zookeeper 0.842 4.429 2.280 8.000 15.111 24.192
Average 0.694 12.554 3.091 17.831 29.870 41.552
Summary for RQ2: LogPrism achieves state-of-the-art end-to-end compression speed. This efficiency stems from the synergistic effect of its hierarchical processing strategy, which minimizes computational overhead, and its parallel-aware architecture, which maximizes resource utilization.
IV-D RQ3: Ablation Analysis of LogPrism’s Different Stages
We perform a progressive ablation study to isolate and quantify the contribution of each stage in LogPrism’s hierarchical design. Specifically, we focus on Stage 2 (Variable Subtree Encoding), which implements our core unified redundancy encoding paradigm. We evaluate three configurations across all 16 datasets: (1) LZMA, a general-purpose compressor serving as a universal baseline; (2) LogPrism (S1+S3), a structural baseline that executes Stage 1 and Stage 3 but bypasses Stage 2, representing an advanced parser-based compressor that treats structure and variables as separate entities; and (3) LogPrism, the complete model, which is compared against LogPrism (S1+S3) to measure the specific performance gain provided by mining “structure + variable” correlations. Both the last two configurations utilize LZMA for the final compression of their output. The results for CR and CS are detailed in Tables IV and V, respectively.
[…]
TABLE IV: Ablation Study of Compression Ratio
Dataset LZMA LogPrism (S1+S3) LogPrism
Android 18.857 25.564 32.822
Apache 25.186 61.887 64.330
BGL 17.637 46.548 47.145
Hadoop 36.095 75.170 79.733
HDFS 13.559 24.282 26.466
HealthApp 13.431 46.932 50.477
HPC 15.076 43.880 45.447
Linux 16.677 37.356 38.343
Mac 22.159 37.373 44.992
OpenSSH 18.918 95.969 113.327
OpenStack 14.437 21.868 23.356
Proxifier 18.982 26.664 29.849
Spark 19.908 50.656 54.813
Thunderbird 27.309 59.876 75.862
Windows 202.568 490.787 497.298
Zookeeper 27.667 118.531 143.742
Compression Speed Analysis: The ablation study also reveals the critical role of Stage 2 as a performance accelerator. Despite adding another analysis step, the full LogPrism model is $25.93\%$ faster on average than the structurally simpler LogPrism (S1+S3) variant. This counter-intuitive result validates our hierarchical redundancy mining strategy. In the LogPrism (S1+S3) configuration, every variable identified in Stage 1 must be processed by Stage 3, which relies on computationally expensive regex templatization. In contrast, the full LogPrism model only supplies Stage 3 with a smaller set of “true long-tail residuals.” By handling the majority of high-frequency variables in the efficient Stage 2, LogPrism drastically reduces the workload of the most time-consuming part of the pipeline, thereby lowering overall computational overhead and increasing throughput.
TABLE V: Ablation Study of Compression Speed (MB/s)
Dataset LZMA LogPrism (S1+S3) LogPrism
Android 24.876 26.717 36.672
Apache 7.665 13.453 17.551
BGL 16.108 49.072 50.625
Hadoop 22.102 44.937 52.840
HDFS 15.715 31.145 35.288
HealthApp 9.436 23.156 29.103
HPC 11.700 33.401 37.868
Linux 4.592 8.818 10.228
Mac 8.428 12.950 16.097
OpenSSH 15.873 46.553 70.016
OpenStack 11.647 20.578 29.396
Proxifier 6.696 8.750 10.585
Spark 20.757 50.734 59.005
Thunderbird 26.845 51.360 69.213
Windows 63.706 91.402 116.152
Zookeeper 6.581 14.907 24.192
Average 17.045 32.996 41.552
Summary for RQ3: The ablation study confirms that LogPrism’s performance breakthrough is driven by its hierarchical design. While Stages 1 and 3 provide a solid structural baseline, the core innovation (i.e., Stage 2 Variable Subtree Encoding) delivers a decisive performance boost to both compression ratio and speed.
IV-E RQ4: Robustness Analysis: The Impact of Data Granularity
This question investigate the impact of data granularity on LogPrism’s performance, which is a critical operational parameter. Our primary experiments utilized a fixed chunk size of $100\text{K}$ lines, a standard baseline setting that maximizes parallelism but limits pattern discovery to local windows. This design involves an inherent trade-off: each chunk builds an independent URT, preventing the sharing of patterns across the full dataset. This can potentially obscure very long-range redundancies. To explore this speed-vs-compression trade-off, we configure LogPrism to operate in single-archive (non-chunked) mode, allowing it to construct a single global URT over the entire dataset. We compare this configuration against Denum, the leading state-of-the-art baseline that also supports global operation. The results, presented in Table VI, demonstrate that global pattern discovery significantly boosts LogPrism’s compression effectiveness while maintaining highly competitive speeds.
Compression Ratio Analysis: In single-archive mode, LogPrism achieves a 273.27% increase in average CR over its own chunked performance, with the gain on the Windows dataset reaching a massive 739.37%. When compared to Denum operating in the same global configuration, LogPrism outperforms it on 13 of the 16 datasets. The improvements are particularly significant on complex logs such as Windows (+22.49%) and Thunderbird (+21.29%). On two of the remaining three datasets (Hadoop, HDFS), LogPrism remains highly competitive, following Denum with a small difference of less than 2.50%. The only notable exception is the Spark dataset, where Denum’s specialized handling of specific numeric sequences proves more efficient. Overall, these results validate that LogPrism scales effectively to leverage global context to achieve superior compression.
[…]
TABLE VI: Robustness Analysis of Single-Archive Mode Performance
Dataset Compression Ratio Speed (MB/s)
Denum LogPrism Denum LogPrism
Android 46.191 52.429 5.416 15.294
Apache 58.562 64.269 5.810 12.369
BGL 44.284 47.429 6.100 16.345
Hadoop 92.899 90.590 8.643 27.186
HDFS 32.919 32.198 5.230 7.956
HealthApp 47.279 53.574 7.072 12.728
HPC 45.542 46.350 6.369 16.276
Linux 30.688 38.216 4.582 8.456
Mac 43.633 50.266 5.245 11.171
OpenSSH 106.984 120.116 7.426 27.962
OpenStack 21.898 23.394 5.129 12.576
Proxifier 27.141 29.739 5.484 9.894
Spark 65.125 56.196 5.398 13.237
Thunderbird 70.161 85.097 4.263 15.900
Windows 3407.766 4174.176 6.421 26.897
Zookeeper 135.846 142.290 7.556 18.013
Average 267.307 319.146 6.009 15.766
Summary for RQ4: LogPrism’s design offers both robustness and flexibility, providing users with a clear speed-vs-compression trade-off: the default chunked mode maximizes speed, while the single-archive mode maximizes the compression ratio at a reasonable efficiency cost. In both operational paradigms, LogPrism’s performance advantages significantly surpass baseline models.
V Related Work
This section discusses the landscape of compression techniques relevant to log data, organized by their foundational design philosophies.
V-A General-purpose Compression Approaches
These algorithms eliminate statistical redundancy in data for compression, which can be categorized into three primary families. Dictionary-based approaches, exemplified by LZMA, achieve compression by replacing repeated byte sequences with references to a dictionary. Prediction-based methods, such as PPMd, leverage statistical models to encode characters based on their preceding context. Block-sorting algorithms, notably bzip2, utilize the Burrows-Wheeler Transform to cluster similar characters, thereby enhancing subsequent encoding efficiency. While these methods exhibit strong performance on generic text data, they fundamentally operate on undifferentiated byte streams without awareness of log structure. Consequently, their effectiveness diminishes significantly when confronted with logs containing high-entropy variables such as unique request IDs and dynamic numerical values, as they cannot exploit the inherent semi-structured redundancy that spans log entries.
V-B Log-specific Compression Approaches
The dominant methodology in log compression is the parser-based paradigm [23, 33, 21, 29, 32], which decouples structure extraction from variable encoding. LogZip established this framework by combining the Drain parser with iterative clustering. Subsequent works focused on optimizing the separated components: LogReducer introduced delta encoding for timestamps, while LogShrink utilized entropy-based analysis to identify variable patterns. Additionally, systems like CLP and LogGrep have extended this paradigm to enable direct search on compressed data. To avoid the overhead of explicit template extraction, alternative approaches have emerged. Early works, such as LogArchive [7] and MLC [12], applied similarity grouping or block-level deduplication to compress logs. More recently, the number-centric paradigm, exemplified by Denum [38], performs a global binary classification to split tokens into numeric and non-numeric streams.
However, state-of-the-art methods share a fundamental limitation: the reliance on early, rigid token categorization. Parser-based methods separate templates from variables, while number-centric methods separate numerics from strings. This irreversible decoupling destroys contextual correlations that span these boundaries. Instead, LogPrism unifies structural extraction and pattern encoding to mine deep “structure+variable” redundancies that existing methods ignore.
VI Conclusion
[…]
Data Availability
The source code of LogPrism is publicly available on https://github.com/Lycc42/LogPrism.